given that AUB=147 A=81, and B=66 find ANB.
Answer by user_dude2008(1862)

n(AUB)=n(A)+n(B)-n(A N B)


147=81+66-n(A N B)

147=147-n(A N B)

0=-n(A N B)


Ans: n(A N B) = 0